OpenStudy (anonymous):

Determine the 2 x-intercepts and the vertex of a parabola in (x,y) form and how you found these ordered pairs in a sentence. (-1,0) (2,0)

OpenStudy (anonymous):

I am not clear on the question. Also, what is the equation?

OpenStudy (anonymous):

the first question was to find the corresponding y values when x= -2,-1,0,1,2,3 for the equation y=x^2-x-2

OpenStudy (anonymous):

i got (-2,4) (-1,0) (0,-2) (1,-2) (2,0) and (3,4)

OpenStudy (anonymous):

To find x-intercepts, Let y = 0 and solve for x. So: \[y=x^2-x-2 \rightarrow 0=x^2-x-2\] Now you have to factor to solve. (Or quadratic equation) You will get 2 solutions and they want you to write it in (x, y) form. Since you let y=0, both will have 0 for y.

OpenStudy (anonymous):

for the vertex, you use that -b/2a again to find the x-coordinate and then plug it in to find the y-coordinate - like the last problem

OpenStudy (anonymous):

so (x-1)(x+2)

OpenStudy (anonymous):

and the vertex would be -1/2

OpenStudy (anonymous):

You are close. \[\left( x-1 \right)\left( x+2 \right)=x ^{2}+x-2\] but you want \[x ^{2}-x-2\] so your signs are off.

OpenStudy (anonymous):

so it would be (x-1)(x-2)

OpenStudy (anonymous):

no (x+1)(x-2)

OpenStudy (anonymous):

You were right that you needed opposite signs

OpenStudy (anonymous):

yes ... now set each factor = 0 and solve each equation.

OpenStudy (anonymous):

but when I graph my (x,y) from the first question. the graph shows the x-intercepts as (-1,0) (2,0)

OpenStudy (anonymous):

when you set x + 1 = 0 and solve, - 1 - 1 We get: x = 0

OpenStudy (anonymous):

What do you get when you set x - 2 = 0?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yes. Now remember you set y = 0 and solved for x, so the points you found are (1, 0) and (-2, 0)

OpenStudy (anonymous):

Now, back to the vertex. When you set x = -b/2a, you found the x-coordinate of the vertex to be 1/2. What is the y-coordinate?

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

\[y=x ^{2}-x-2\] so \[y=\left( \frac{ 1 }{ 2 } \right)^{2}-\frac{ 1 }{ 2 }-2\]

OpenStudy (anonymous):

I think maybe you made a sign error. \[\left( \frac{ 1 }{ 2 } \right)^{2}=\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

so: \[\frac{ 1 }{ 4 }-\frac{ 1 }{ 2 }-2 = ?\]

OpenStudy (anonymous):

i got -1/2 for my vertex

OpenStudy (anonymous):

only for the x-coordinate. The vertex is a point. You need the y-coordinate. You have to replace x with 1/2 to find the corresponding y.

OpenStudy (anonymous):

Try this. 1/4 = 0.25 and 1/2 = 0.5, so we have 0.25-0.5-2

OpenStudy (anonymous):

-2.25

OpenStudy (anonymous):

yes! Or - 2 1/4

OpenStudy (anonymous):

Now express it as a point, so (1/2, -2.25)

OpenStudy (anonymous):

And it sounded like they want you to explain in words what you did.

OpenStudy (anonymous):

In case it might be useful to you later, this is a list of my videos for factoring and solving quadratic equations by factoring: www.mathmods.com/mat100/Chapters/Factoring/factoring_video_links.pdf

OpenStudy (anonymous):

@davisla I am still having trouble with this problem. I figured out the factoring and the vertex but when I graph this from the first part the x-intercepts that I get from graphing are (-1,0) (2,0). Can you please help me?

OpenStudy (anonymous):

I will make you a video of the whole problem in about 15 min if that is ok

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

here is the video ... let me know if this doesn't do it for you: www.mathmods.com/Math_Tutoring/vertex_int_parabola_5_5_min/vertex_int_parabola_5_5_min.html

OpenStudy (anonymous):

Sorry, I didn't read your question correctly ... here is the graph: www.mathmods.com/Math_Tutoring/vertex_int_parabola_1_min/vertex_int_parabola_1_min.html